OpenStudy (anonymous):

Write in point-slope form an equation of the line through each pair of points.

OpenStudy (anonymous):

(-10,3) and (-2,-5)

OpenStudy (anonymous):

@hartnn @Haseeb96 i'm not really sure what these directions mean

OpenStudy (anonymous):

slope of -1

OpenStudy (anonymous):

equation is y+5=-1(x+2)

OpenStudy (anonymous):

is that all?

OpenStudy (anonymous):

y=-x+7

OpenStudy (anonymous):

someone check me

OpenStudy (haseeb96):

First find slope by using this formula m=y2-y1/x2-x1 now just put the above points in this formula and get the slope first and show me your work.

OpenStudy (anonymous):

oh okay so that's how he determined the slope...

OpenStudy (anonymous):

was he right haseeb?

OpenStudy (anonymous):

prob not since i did it in my head

OpenStudy (haseeb96):

Yes the slope is -1 Now the formula of point slope form is (y-y1)=m(x-x1) Now just put this slope and any one point from the above two points and get the answer.

OpenStudy (anonymous):

this is kind of confusing lol.

OpenStudy (haseeb96):

The correct answer is y=-x-7

OpenStudy (anonymous):

oh..okay